Consider the following distributions for a random variable X. In each case, match the distribution with the expectation of Y= 10-2x2 Poisson distribution with ?=5-50 Binomial(n= 10,p=0.3) Geometric(p-0.6) 12.2 4 0 x-oo! Choose.. ? Just the last part please :)
Answer #1

E(X^2) = \int_{0}^{\infty }x^2 *f(x)dx = \int_{0}^{\infty }x^2 * \frac{1}{2}e^{-\frac{x}{2}}dx

\Rightarrow E(X^2) = \int_{0}^{\infty }x^2 * \frac{1}{2}e^{-\frac{x}{2}}dx

\Rightarrow E(X^2) = \frac{1}{2}\int_{0}^{\infty }x^2 *e^{-\frac{x}{2}}dx

\Rightarrow E(X^2) = \frac{1}{2} * \frac{\Gamma (3)}{(\frac{1}{2})^3}

\Rightarrow E(X^2) = \frac{1}{2} * \frac{2!}{(\frac{1}{2})^3}

\Rightarrow E(X^2) = \frac{2!}{(\frac{1}{2})^2}

\Rightarrow E(X^2) = 2*2^2 = 8

Now,

E(Y)= E(10-2X^2)= 10-2E(X^2) = 10 - (2*8) = 10 - 16 = -6